### v2

# Gradient as Conditions: Rethinking HOG for All-in-one Image Restoration

Abstract
All-in-one image restoration (AIR) aims to address diverse degradations within a unified model by leveraging informative degradation conditions to guide the restoration process. However, existing methods often rely on implicitly learned priors, which may entangle feature representations and hinder performance in complex or unseen scenarios. Histogram of Oriented Gradients (HOG) as a classical gradient representation, we observe that it has strong discriminative capability across diverse degradations, making it a powerful and interpretable prior for AIR. Based on this insight, we propose HOGformer, a Transformer-based model that integrates learnable HOG features for degradation-aware restoration. The core of HOGformer is a Dynamic HOG-aware Self-Attention (DHOGSA) mechanism, which adaptively models long-range spatial dependencies conditioned on degradation-specific cues encoded by HOG descriptors. To further adapt the heterogeneity of degradations in AIR, we propose a Dynamic Interaction Feed-Forward (DIFF) module that facilitates channel–spatial interactions, enabling robust feature transformation under diverse degradations. Besides, we propose a HOG loss to explicitly enhance structural fidelity and edge sharpness. Extensive experiments on a variety of benchmarks, including adverse weather and natural degradations, demonstrate that HOGformer achieves state-of-the-art performance and generalizes well to complex real-world scenarios. Code is available at https://github.com/Fire-friend/HOGformer.
Introduction
Figure 1: Visualization of HOG feature distributions under various degradations. (a) Example images of different weather conditions with corresponding HOG feature visualizations. (b) HOG features for five natural degradations (Zheng et al. 2024) and three adverse weather degradations (Sun et al. 2024), using 100 randomly selected images for each degradation.
Real-world images are frequently affected by diverse degradations such as blur, low-light, and adverse weather. While task-specific deep learning models have achieved remarkable progress in restoring images under individual degradations (He and Jin 2024; Wang, Wu, and Jin 2023; Jin et al. 2025), they often require separate models, leading to high training and deployment overhead. AIR (Jiang et al. 2024a) offers a practical alternative by handling diverse degradations in a single model, which is critical for real-world applications like autonomous driving (Chen et al. 2024).
A key challenge in AIR is designing a model that can adaptively respond to the specific degradation present in each input. To address this, recent works (Li et al. 2022; Potlapalli et al. 2023; Ai et al. 2024; Jiang et al. 2024b) have proposed shared backbone architectures modulated by degradation-aware conditional features, enabling a unified framework to handle diverse tasks. Among them, prompt-based methods such as PromptIR (Potlapalli et al. 2023) inject degradation-specific prompts to guide the network behavior, while others exploit multimodal inputs (Ai et al. 2024) to capture complementary degradation cues. Despite their promise, the success of these methods critically depends on the quality and reliability of the conditioning. In practice, such signals are often derived from implicitly learned priors, which operate as black boxes with limited interpretability. This lack of transparency makes it difficult to disentangle the underlying degradations, especially when they exhibit subtle or overlapping characteristics. Consequently, feature representations optimized for one degradation (e.g., denoising) may conflict with those required for another (e.g., dehazing), leading to feature entanglement and reduced generalization performance across tasks (Li et al. 2020). While alternative methods based on Mixture-of-Experts (MoE) (Zamfir et al. 2025; Ai, Huang, and He 2024) offer greater flexibility through expert routing, they come with substantial computational overhead, making them less suitable for efficient deployment. These limitations underscore the requirement for a conditional mechanism that is not only explicit and discriminative, but also computationally efficient and robust to diverse degradations.
In this work, we revisit classical feature descriptors and identify the HOG (Dalal and Triggs 2005) features as a surprisingly effective and interpretable prior for AIR. As shown in Figure 1, we observe that the gradient patterns captured by HOG form distinct signatures for different degradations. For example, in the case of rain (Figure 1(a)), falling rain produces vertical streaks that are reflected as strong vertical gradients, while raindrop regions appear as isolated circular patches with low gradient magnitudes. In contrast, snow degradation introduces widespread high-magnitude gradients with low directional variance, forming dense and uniform textures. This discriminative behavior persists across both natural and adverse weather conditions (Figure 1(b)). The effectiveness of HOG stems from its two complementary components, i.e., gradient magnitude and orientation, which explicitly encode local intensity variations and directional structure. These observations raise a central question: Can the inherent degradation-discriminative property of HOG be leveraged to guide the design of AIR networks, enabling them to explicitly capture degradation-specific gradient patterns for more efficient AIR?
To this end, we propose HOGformer, an AIR framework that integrates HOG as gradient priors in a principled and degradation-aware manner. Instead of incorporating the static handcrafted HOG feature, we reformulate it as a learnable and differentiable module that provides dynamic and context-aware guidance throughout the network. At the core of HOGformer lies the Dynamic HOG-aware Self-Attention (DHOGSA) mechanism, which integrates HOG-derived cues into the self-attention process. This design enables the network to focus on degradation-specific patterns, such as rain streaks or haze-induced texture loss, by explicitly modulating attention maps based on gradient information. To further enhance the adaptability of feature representations, we introduce a Dynamic Interaction Feed-Forward (DIFF) module that facilitates effective channel–spatial interaction, allowing the model to robustly transform features under diverse and heterogeneous degradations. In addition, we propose a dedicated HOG supervision loss that directly constrains the reconstruction of gradient magnitude and orientation. This explicit guidance promotes sharper edges and better preservation of structural details. This unified design enables effective restoration across diverse degradations, achieving state-of-the-art (SOTA) performance on both adverse weather and natural image AIR tasks. Our main contributions are summarized as follows:
  • We identify that classical HOG features serve as an explicit and highly discriminative prior for distinguishing diverse degradations in AIR, offering a compelling alternative to implicit conditional mechanisms.
  • We propose HOGformer, an AIR network that embeds learnable HOG cues into core network components for degradation-aware adaptation. HOGformer integrates a dynamic HOG-aware self-attention (DHOGSA) mechanism, an efficient dynamic interaction feed-forward (DIFF) module, and a dedicated HOG supervision loss.
  • Extensive experiments show that HOGformer achieves state-of-the-art performance across various degradations, including adverse weather and natural conditions.
Figure 2: The overall architecture of our HOGformer. It includes the HOG Transformer block with the Dynamic HOG-aware Self-Attention (DHOGSA) module and Dynamic Interaction Feed-Forward (DIFF) module.
Related Works
Single-Task Image Restoration
Image restoration is a fundamental problem in computer vision. Traditional methods constrain the solution space using human priors and handcrafted features (Banham and Katsaggelos 1997). With deep learning, various approaches have achieved strong performance across tasks (He and Jin 2024; Wang, Wu, and Jin 2023; Qiu et al. 2023; Zamir et al. 2022a; Guo et al. 2024). Vision Transformers further improve restoration by modeling long-range dependencies (Liang et al. 2021; Zamir et al. 2022a; Wang et al. 2022a). For example, Swin-IR (Liang et al. 2021) adopts a shifted window strategy for local-global context modeling, while Restormer (Zamir et al. 2022a) introduces a multi-stage attention design to balance accuracy and efficiency. Transformer-based models have also been applied to deraining (Chen et al. 2023), desnowing (Chen et al. 2022b), dehazing (Song et al. 2023; Zhang et al. 2024), deblurring (Liang et al. 2024; Kong et al. 2023), and low-light enhancement (Cai et al. 2023; Zhang, Meng, and Lam 2023). However, these methods require separate training, increasing computational and deployment costs (Sun et al. 2024).
All-in-One Image Restoration
AIR aims to handle diverse degradations using a single model without task-specific retraining (Jiang et al. 2024a; Li et al. 2022; Sun et al. 2024; Zheng et al. 2024). Typically, InstructIR (Conde, Geigle, and Timofte 2024) employs natural language instructions to specify restoration goals but incurs high data preparation costs. Methods such as Painter (Wang et al. 2023) and DA-CLIP (Luo et al. 2024) leverage on-the-fly learning to adapt large models, while DiffUIR (Zheng et al. 2024) builds on residual diffusion to address diverse degradations. PromptIR (Potlapalli et al. 2023) encodes degradation conditions into prompts to guide restoration. These approaches rely on implicitly constructed degradation-specific conditions to enable flexible inference. However, the construction of such conditions is often complex and lacks consistent priors. Recently, Histoformer (Sun et al. 2024) incorporates grayscale histogram priors into Transformer models and shows strong performance on adverse weather restoration. However, histogram-based features offer limited discriminative power; for example, blurry and clean images can share similar histograms. In this work, we observe that HOG offers a clearer way to distinguish between different degradations. Our method utilizes HOG features to provide explicit guidance that is aware of the degradation throughout the restoration process.
Method
Overall Architecture
Motivated by the observation (Figure 1) that the HOG effectively captures degradation-specific patterns, we propose HOGformer, an AIR network that explicitly incorporates HOG-derived cues into the restoration process. As shown in Figure 2, HOGformer employs a multi-stage U-shaped encoder-decoder architecture to progressively model and restore diverse degradations within a single model.
The encoder begins by applying a $3\times 3$ convolution to extract shallow features from the input image $I^{lq}$. These features are successively downsampled and enriched through stacked HOG Transformer Blocks (HOGTBs), which serve as the core feature extractors. To retain low-frequency priors and facilitate residual learning, coarse skip connections (Sun et al. 2024) are introduced to highlight degradation residuals, incorporating average pooling, pointwise convolution, and depthwise convolution. Downsampling and upsampling are achieved using pixel-unshuffle and pixel-shuffle, respectively, which ensure efficient resolution transitions while preserving structural consistency. Additionally, symmetric skip connections between the encoder and decoder further enhance information flow and detail recovery. The core of the HOGformer lies the HOGTBs, which incorporate dynamic HOG-aware self-attention (DHOGSA) and dynamic interaction feed-forward (DIFF) as illustrated in Figure 2. DHOGSA utilizes gradient information to adjust attention through HOG-guided sorting at both pixel- and patch-level. This facilitates selective focus on regions sensitive to degradation regions. DIFF facilitates effective channel–spatial interaction to enable robust transformation of features under diverse degradations. Each HOGTB employs a residual paradigm to ensure stable optimization as follows:
$\begin{split}\mathbf{F}^{\prime}_{l}&=\mathbf{F}_{l-1}+\text{DHOGSA}(\text{LN}(\mathbf{F}_{l-1})),\\ \mathbf{F}_{l}&=\mathbf{F}^{\prime}_{l}+\text{DIFF}(\text{LN}(\mathbf{F}^{\prime}_{l})),\end{split}$ (1)
where $F_{l}$ denotes the output of the $l$-th layer, LN is the layer normalization. This structured integration of HOG-guided attention and interaction enables efficient AIR.
Dynamic HOG-aware Self-Attention (DHOGSA)
The first core module in each HOGTB is DHOGSA, designed to capture long-range and degradation-specific dependencies using gradient-domain priors. Traditional self-attention mechanisms (Zamir et al. 2022a) typically operate within fixed windows or channel-wise structures, making them suboptimal for modeling the non-uniform and spatially varying patterns present in diverse degradations. DHOGSA overcomes this by leveraging differentiable HOG descriptors to explicitly sort features based on their gradient magnitudes and orientations before computing attention. This sorting operation is conducted at both patch and pixel levels, grouping spatial locations with similar degradation patterns to facilitate more effective attention computation.
Local Dynamic-range Convolution (LDRConv).
To facilitate effective long-range modeling, we begin by enhancing local degradation structures through a Local Dynamic-Range Convolution (LDRConv) module. Unlike standard convolutions with fixed receptive fields, LDRConv dynamically reorganizes features based on their gradient distribution. Specifically, LDRConv performs HOG-guided patch-level sorting, followed by modulation using learnable bin-wise HOG priors. This patch-wise operation groups regions with similar gradient properties while preserving global spatial consistency, in contrast to pixel-wise sorting which may disrupt the overall structure. This design ensures that the model retains critical semantic and geometric information, which is essential for image restoration tasks.
Given input features $\mathbf{F}\in\mathbb{R}^{C\times H\times W}$, we first compute the gradient magnitude $m$ and orientation $o$ using Sobel filters:
$\mathbf{m}=\sqrt{\mathbf{g_{x}}^{2}+\mathbf{g_{y}}^{2}},o=\left\lfloor\frac{\operatorname{atan2}(\mathbf{g_{y}},\mathbf{g_{x}})+\pi}{2\pi}N_{\text{bin}}\right\rfloor,$ (2)
where $\mathbf{g_{x}}$ and $\mathbf{g_{y}}$ are obtained directional gradients using Sobel operation, $N_{\text{bin}}$ is the number of HOG bins. To preserving the original representation capacity, we split $\mathbf{F}$ into $\mathbf{F}_{1},\mathbf{F}_{2}$ and only apply HOG-guided patch-wise sorting (Figure 3(a)) to $\mathbf{F}_{1}$, followed by modulation with learnable HOG priors $\operatorname{HOG}_{\theta}(\mathbf{F}_{1})$ to enhance local gradient awareness under diverse degradations:
$\begin{split}&\mathbf{F}_{1}=\operatorname{Sort_{patch}}(\mathbf{F}_{1})+\operatorname{HOG}_{\theta}(\mathbf{F}_{1}),\\ &\mathbf{F}=\operatorname{Conv_{3\times 3}^{d}}(\operatorname{Conv_{1\times 1}^{p}}(\operatorname{Concat}(\mathbf{F}_{1},\mathbf{F}_{2}))).\end{split}$ (3)
As shown in Figure 3(c), $\operatorname{HOG}_{\theta}(\mathbf{F}_{1})$ is derived by computing gradient-based histograms over local patches, followed by feature projection, more details can be found in the supplementary material. LDRConv enhances degradation sensitivity while maintaining structural coherence, providing a strong basis for the subsequent self-attention mechanism.
Figure 3: HOG-guided mechanism: (a) Patch-level sorting. (b) Pixel-level sorting. (c) HOG feature extraction process.
HOG-guided Self-Attention.
With the enhanced representations from LDRConv, we effectively model the crucial long-range dependencies. Specifically, inspired by the observation that distant pixels affected by the same degradation often exhibit similar HOG responses, we sort pixels (Figure 3(b)) based on HOG descriptors ($\mathbf{m}\cdot\mathbf{o}$). To spatially align features, we generate sorting indices from the information-rich value features $\mathbf{V}$, and then use these indices to consistently sort the query ($\mathbf{Q}$), key ($\mathbf{K}$), and value ($\mathbf{V}$) features. To capture multi-scale dependencies, we introduce two parallel attention branches with distinct histogram reshaping strategies. Bin-wise Histogram Reshaping (BHOGR) groups sorted pixels into fixed-size bins to extract coarse, large-scale features. Frequency-wise Histogram Reshaping (FHOGR) clusters pixels with similar HOG values to capture fine textures and local variations. BHOGR focuses on capturing large-scale structural artifacts, such as the spatially uniform distribution of haze, whereas FHOGR is tailored to detect fine-scale, repetitive degradations like rain streaks. Fusing both branches enables a comprehensive representation $\mathbf{F}$. The entire process is as follows:
$\begin{split}&\mathbf{idx}=\operatorname{Sort}\Bigl{(}\mathcal{R}_{C\times H\times W}^{C\times HW}(o(\mathbf{V})\cdot m(\mathbf{V}))\Bigr{)},\\ &\mathbf{Q}_{\{\cdot\}\in\{B,F\}}=\operatorname{Gather}\Bigl{(}\mathcal{R}_{C\times H\times W}^{C\times HW}(\mathbf{Q}_{\{\cdot\}}),\,\mathbf{idx}\Bigr{)},\\ &\mathbf{K}_{\{\cdot\}\in\{B,F\}}=\operatorname{Gather}\Bigl{(}\mathcal{R}_{C\times H\times W}^{C\times HW}(\mathbf{K}_{\{\cdot\}}),\,\mathbf{idx}\Bigr{)},\\ &\mathbf{V}=\operatorname{Gather}\Bigl{(}\mathcal{R}_{C\times H\times W}^{C\times HW}(\mathbf{V}),\,\mathbf{idx}\Bigr{)},\\ &\mathbf{A}_{\{\cdot\}\in\{B,F\}}=\text{Softmax}\left(\frac{\mathcal{R}_{\{\cdot\}}(Q_{\{\cdot\}})\mathcal{R}_{\{\cdot\}}(K_{\{\cdot\}})^{\top}}{\sqrt{\mathbf{K}}}\right),\\ &\mathbf{F}=\mathbf{A_{B}}\mathcal{R}_{B}(\mathbf{V})\odot\mathbf{A_{F}}\mathcal{R}_{F}(\mathbf{V}).\end{split}$ (4)
where $\mathbf{K}$ is the number of heads, $\mathcal{R}_{i\in\{B,F\}}$ represents the reshaping operation (BHOGR or FHOGR), and $\odot$ is the Hadamard product.
Snow100K-S Snow100K-L Outdoor-Rain RainDrop Average
Method P $\uparrow$ S $\uparrow$ P $\uparrow$ S $\uparrow$ Method P $\uparrow$ S $\uparrow$ Method P $\uparrow$ S $\uparrow$ P $\uparrow$ S $\uparrow$
RESCAN 31.51 0.9032 26.08 0.8108 HRGAN 21.56 0.8550 RaindropAttn 31.44 0.9263 - -
SnowGAN 32.33 0.9500 27.17 0.8983 PCNet 26.19 0.9015 AttentiveGAN 31.59 0.9274 - -
DDMSNet 34.34 0.9445 28.85 0.8772 MPRNet 28.03 0.9192 IDT 31.87 0.9313 - -
DTANet 34.79 0.9497 30.06 0.9017 NAFNet 29.59 0.9027 MAXIM 31.87 0.9352 - -
Restormer 36.01 0.9579 30.36 0.9068 Restormer 30.03 0.9215 Restormer 32.18 0.9408 - -
All-in-One - - 28.33 0.8820 All-in-One 24.71 0.8980 All-in-One 31.12 0.9268 - -
TransWeather 32.51 0.9341 29.31 0.8879 TransWeather 28.83 0.9000 TransWeather 30.41 0.9157 30.27 0.9094
Chen et al. 34.42 0.9469 30.22 0.9071 Chen et al. 29.27 0.9147 Chen et al. 31.81 0.9309 31.43 0.9249
WGWSNet 34.31 0.9460 30.16 0.9007 WGWSNet 29.32 0.9207 WGWSNet 32.38 0.9378 31.54 0.9263
WeatherDiff64 35.83 0.9566 30.09 0.9041 WeatherDiff64 29.64 0.9312 WeatherDiff64 30.71 0.9312 31.57 0.9308
WeatherDiff128 35.02 0.9516 29.58 0.8941 WeatherDiff128 29.72 0.9216 WeatherDiff128 29.66 0.9225 30.99 0.9225
AWRCP 36.92 0.9652 31.92 0.9344 AWRCP 31.39 0.9329 AWRCP 31.93 0.9314 33.04 0.9409
GridFormer 37.46 0.9640 31.71 0.9231 GridFormer 31.87 0.9335 GridFormer 32.39 0.9362 33.36 0.9392
MoCE-IR 37.10 0.9654 31.88 0.9234 MoCE-IR 31.42 0.9334 MoCE-IR 32.23 0.9386 33.16 0.9400
Histoformer 37.41 0.9656 32.16 0.9261 Histoformer 32.08 0.9389 Histoformer 33.06 0.9441 33.67 0.9436
HOGformer-L 37.93 0.9685 32.41 0.9297 HOGformer-L 32.89 0.9460 HOGformer-L 32.72 0.9452 33.99 0.9474
Table 1: Quantitative comparisons of adverse weather restoration. The top half of the tables shows results from task-specific methods, while the bottom half displays evaluations of AIR methods. The best and second-best results are highlighted.
Figure 4: Visual comparison for deraining (Sun et al. 2024; Özdenizci and Legenstein 2023). Zoom in for the best visualization.
Rain (5 sets) Low-light Snow (2 sets) Haze Blur Complexity
Method P $\uparrow$ S $\uparrow$ P $\uparrow$ S $\uparrow$ P $\uparrow$ S $\uparrow$ P $\uparrow$ S $\uparrow$ P $\uparrow$ S $\uparrow$ Params (M) FLOPs (G)
SwinIR 30.78 0.923 17.81 0.723 - - 21.5 0.891 24.52 0.773 0.90 752.13
MIRNet-v2 33.89 0.924 24.74 0.851 - - 24.03 0.927 26.30 0.799 5.90 140.92
Restormer 33.96 0.935 20.41 0.806 - - 30.87 0.969 32.92 0.961 26.12 141.00
MAXIM 33.24 0.933 23.43 0.863 - - 34.19 0.985 32.86 0.940 14.1 216.00
IR-SDE - - - - 20.45 0.787 - - 30.70 0.901 34.20 98.30
WeatherDiff - - - - 33.51 0.939 - - - - 82.96 -
RDDM 30.74 0.903 23.22 0.899 32.55 0.927 30.78 0.953 29.53 0.876 36.26 9.88
AirNet 24.87 0.773 14.83 0.767 27.63 0.860 25.47 0.923 26.92 0.811 8.93 30.13
Painter 29.49 0.868 22.40 0.872 - - - - - - 307.00 248.90
ProRes 30.67 0.891 22.73 0.877 - - 32.02 0.952 27.53 0.851 307.00 248.90
Prompt-IR 29.56 0.888 22.89 0.847 - - 32.02 0.952 27.21 0.817 35.59 15.81
DA-CLIP 28.96 0.853 24.17 0.882 30.80 0.888 31.39 0.983 25.39 0.805 174.10 118.50
DiffUIR-S 30.25 0.893 23.52 0.895 31.45 0.915 31.83 0.954 27.79 0.830 3.27 2.40
DiffUIR-L 31.03 0.904 25.12 0.907 32.65 0.927 32.94 0.956 29.17 0.864 36.26 9.88
HOGformer-S 30.75 0.901 25.36 0.915 32.72 0.929 33.67 0.991 28.37 0.840 2.91 20.63
HOGformer-L 31.63 0.914 25.57 0.917 34.08 0.941 36.60 0.994 29.95 0.884 16.64 91.77
Table 2: Quantitative results of all-in-one image restoration methods in five tasks. The best and second-best results are highlighted. The top is task-specific restoration methods, and the bottom is all-in-one restoration methods.
Dynamic Interaction Feed-Forward (DIFF)
After DHOGSA aggregates spatial features guided by HOG cues, the resulting representations are fed into the proposed Dynamic Interaction Feed-Forward (DIFF) network. Unlike standard, content-agnostic FFNs, DIFF performs dynamic spatial-channel interaction to adaptively refine features based on the content, so that to enhance heavily degraded regions while preserving clean areas. Three complementary designs make it possible: firstly, multi-scale spatial context via parallel branches is captured; secondly, a pixel-wise gating mechanism modulates spatial responses conditioned on input features; thirdly, channel shuffling and aggregation promote cross-channel communication. Therefore, DIFF can achieve content-aware refinement across spatial and channel dimensions. Further details are provided in the supplementary material.
Loss Supervision
To fully exploit the HOG-guided attention and dynamic feature interaction in HOGformer, we design a tailored optimization objective. While pixel-wise losses such as the $L_{1}$ loss ensure reconstruction accuracy, they often overlook structural and textural fidelity, resulting in overly smooth outputs. To address this limitation, we formulate a composite loss that aligns with the architectural design:
$\mathcal{L}=\mathcal{L}_{\text{rec}}+\alpha\mathcal{L}_{\text{cor}}+\beta\mathcal{L}_{\text{HOG}},$ (5)
where $\alpha$ and $\beta$ are weighting factors. $\mathcal{L}_{\text{rec}}$ denotes the $L_{1}$ loss for pixel-level accuracy, $\mathcal{L}_{\text{cor}}$ promotes global photometric consistency via Pearson correlation (Sun et al. 2024), and $\mathcal{L}_{\text{HOG}}$ enforces structural alignment by minimizing the $L_{2}$ distance of HOG features between the output and GT. More details are provided in the supplementary material.
Figure 5: Comparisons on multiple degradations (rain-fog-snow). Top: PSNR and SSIM comparisons between Histformer and our method. Bottom: Visual results on a representative example.
Experiments
Experimental Setup
We evaluate HOGformer in two standard AIR settings: (I) 3-task adverse weather removal (Sun et al. 2024), including desnowing, draining&dehazing, and raindrop removal; and (II) 5-task general restoration (Zheng et al. 2024), including deraining, low-light enhancement, desnowing, dehazing, and deblurring. One AIR model is trained for each setting.
Datasets and Metrics. For Setting (I), we use the AllWeather dataset (Sun et al. 2024). For Setting (II), we use task-specific datasets: Merged rain dataset (Jiang et al. 2020; Wang et al. 2022b) for deraining, LOL (Wei et al. 2018) for low-light, Snow100K (Liu et al. 2018) for desnowing, RESIDE (Li et al. 2018a) for dehazing, and GoPro (Nah, Hyun Kim, and Mu Lee 2017) for deblurring. To evaluate generalization, we additionally test on real-world datasets: Practical (Yang et al. 2017), MEF (Ma, Zeng, and Wang 2015), NPE (Wang et al. 2013), DICM (Lee, Lee, and Kim 2013), HIDE (Shen et al. 2019), RealBlur (Rim et al. 2020), and T-OLED (Zhou et al. 2021).
[…]
Implementation Details
We implement our model using PyTorch and conduct experiments on NVIDIA Tesla A100 GPUs. We apply random horizontal and vertical flips for data augmentation. Following Histoformer (Sun et al. 2024), $\alpha$ is set to 1, and B equals the number of attention heads. $N_{bin}$ is set to 9 according to original HOG setting (Dalal and Triggs 2005). More details can be seen in the supplementary material.
Rain Low-light Snow Blur
Method N $\downarrow$ N $\downarrow$ N $\downarrow$ P $\uparrow$ S $\uparrow$
WeatherDiff - - 2.96 - -
CLIP-LIT - 3.70 - - -
Restormer 3.50 3.80 - 32.12 0.926
RDDM 3.34 3.57 2.76 30.74 0.894
AirNet 3.55 3.45 2.75 16.78 0.628
Prompt-IR 3.52 3.31 2.79 22.48 0.770
DA-CLIP 3.42 3.56 2.72 17.51 0.667
DiffUIR-L 3.38 3.14 2.74 30.63 0.890
HOGformer-L 3.31 3.08 2.69 30.92 0.907
Table 3: Quantitative results of known task generalization. The best and second-best results are highlighted. The top is task-specific methods, and the bottom is AIR methods.
Method P $\uparrow$ S $\uparrow$ L $\downarrow$
NAFNet 26.89 0.774 0.346
MPRNet 23.33 0.807 0.383
SwinIR 17.72 0.661 0.519
Restormer 20.98 0.632 0.360
RDDM 17.00 0.626 0.545
AirNet 22.73 0.739 0.374
IDR 27.91 0.793 0.346
Prompt-IR 20.47 0.669 0.462
DA-CLIP 15.74 0.606 0.472
DiffUIR-L 29.55 0.887 0.281
HOGformer-L 29.33 0.889 0.271
Table 4: Quantitative results of unknown task on T-OLED. The best and second-best results are highlighted. The top is task-specific methods, and the bottom is AIR methods.
Comparison with State-of-the-Arts
Setting (I). Table 1 compares HOGformer with task-specific (Li et al. 2018b; Yang et al. 2017; Zhang et al. 2021; Chen et al. 2022a; Zamir et al. 2022a; Li, Cheong, and Tan 2019; Li et al. 2016; Zamir et al. 2021; Chen et al. 2022b; Qian et al. 2018; Quan et al. 2021; Tu et al. 2022) and all-in-one SOTA methods (Li, Tan, and Cheong 2020; Valanarasu, Yasarla, and Patel 2022; Chen et al. 2022c; Zhu et al. 2023; Özdenizci and Legenstein 2023; Ye et al. 2023; Wang et al. 2024; Sun et al. 2024) on synthetic and real-world adverse weather datasets. For fairness, we retrain all-in-one baselines on AllWeather (Li, Tan, and Cheong 2020; Özdenizci and Legenstein 2023). HOGformer consistently outperforms existing methods across all degradations. As shown in Figure 4 and Figure 5, HOGformer also restores clearer details and better handles multiple degradations. This underscores the importance of natural discriminative HOG attributes to model degradations.
Setting (II). Table 2 compares AIR methods under five natural degradations. HOGformer-S achieves competitive results in low-light enhancement, desnowing, and dehazing, outperforming larger models like DA-CLIP (174.1M). Besides, HOGformer-L achieves best results with reasonable parameters than counterparts (Liang et al. 2021; Zamir et al. 2022b, a; Tu et al. 2022; Luo et al. 2023; Özdenizci and Legenstein 2023; Liu et al. 2024; Zamir et al. 2022a; Li et al. 2022; Wang et al. 2023; Ma et al. 2023; Potlapalli et al. 2023; Luo et al. 2024; Zheng et al. 2024). These results highlight its balance between performance and complexity.
Generalization. To further assess generalizability, we evaluate HOGformer on both seen and unseen degradations. As shown in Table 3, HOGformer consistently outperforms previous state-of-the-art methods on seen tasks. In Table 4, HOGformer delivers comparable results on T-OLED with unseen degradations, confirming its robustness and adaptability beyond trained scenarios. This generalization likely stems from the intrinsic properties of HOG features. Unlike learned priors prone to overfitting, HOG captures fundamental structural cues common to many degradations. Therefore, the HOG-guided mechanism can highlight such patterns even in real-world degradations.
LDRConv DHOGSA DIFF $\mathcal{L}_{HOG}$ P $\uparrow$ S $\uparrow$ Params $\downarrow$
$\times$ $\times$ $\times$ $\times$ 30.14 0.9258 14.65M
$\times$ $\times$ $\times$ 30.49 0.9261 14.67M
$\times$ $\times$ 31.68 0.9358 16.77M
$\times$ 32.40 0.9421 16.64M
32.89 0.9460 16.64M
Table 5: Ablation study on the proposed core components.
w/o FHOGR w/o BHOGR DHOGSA
P $\uparrow$ 32.59 31.83 32.89
S $\uparrow$ 0.9432 0.9386 0.9460
Table 6: Ablation study on the FHOGR and BHOGR.
Figure 6: Visualization of the DHOGSA output from the second-to-last HOGTB, highlighting its ability to accurately activate degraded regions.
Ablation Studies
Ablation studies on Outdoor-Rain are conducted to evaluate our methods and to determine optimal hyperparameters.
Effectiveness of Core Components. As shown in Table 5, adding LDRConv significantly improves PSNR and SSIM. This performance gain is attributable to it models local dynamic-range variations commonly found in complex degradations. Introducing DHOGSA further boosts performance by embedding HOG-based priors into the attention mechanism. The reason is that HOG features capture gradient cues that vary across degradations, allowing the model to adjust attention weights adaptively based on perceived degradation as shown in Figure 6. The DIFF module brings additional gains by enabling dynamic feature transformation. The reason is that it enhances the ability of model to process regions with varying degradation levels through spatial-channel interaction. Table 6 shows that both BHOGR and FHOGR are beneficial. We attribute this improvement to BHOGR captures inter-interval dependencies while FHOGR focuses on intra-interval structures. Combining all components yields the best performance, confirming their complementary roles in AIR.
Hyperparameters of HOG Bins. Table 7 analyzes the effect of the number of bins in extracted HOG features. A small number of bins leads to poor performance because it fails to capture essential gradient orientations necessary for identifying degradations. Increasing the number of bins beyond a certain threshold offers only minor improvements while imposing greater computational overhead. To balance accuracy and efficiency, we set the number of bins to 9.
Hyperparameters of Local Dynamic-range Convolution. Figure 7 (a) shows that LDRConv with patch size 8 achieves the best performance (32.89dB PSNR, 0.9460 SSIM), outperforming the baseline by +0.68dB and +0.0067 SSIM. Smaller patches offer limited context, while overly large patches (e.g., size 16) degrade structural fidelity.
Hyperparameters of HOG Loss. As illustrated in Figure 7 (b), setting $\beta=1$ yields the highest performance (32.89dB PSNR, 0.9460 SSIM). A small weight ($\beta=0.1$) offers marginal gains, while an excessive weight ($\beta=10$) hampers performance, indicating the need to balance gradient emphasis and overall fidelity.
13 Bins 9 Bins 5 Bins 1 Bin
P $\uparrow$ 32.93 32.89 32.64 32.32
S $\uparrow$ 0.9464 0.9460 0.9433 0.9417
Table 7: Ablation study on the number of bins in extracted HOG features.
Figure 7: Ablation studies on the patch size of LDRConv (a) and the weight of HOG loss (b).
Conclusion
In this paper, we present HOGformer, an AIR model that leverages the power of HOG features. Through a Dynamic HOG-aware Self-Attention (DHOGSA) mechanism and innovative network components, HOGformer effectively handles diverse image degradations. Our approach demonstrates superior performance while maintaining computational efficiency. Besides, extracting HOG features explicitly helps the model identify various degradations from a gradient perspective and eliminate interfering factors like lighting. This further enhance the generalizability of our method. Future research directions include exploring HOG-based mechanisms in conjunction with emerging architectures (e.g., Mamba) to address the AIR task.